OpenStudy (anonymous):

f(x) = 1/sqrt(x) and g(x) = x^2-5 find g(f(x))

OpenStudy (aum):

\[ g(x) = x^2 - 5 \\ g(f(x)) = (f(x))^2 - 5 \]Plug in f(x) in the above equation and simplify.